OpenStudy (abmon98):

Differentiate with respect to x 1/sin^2x

OpenStudy (kainui):

Do you know how to rewrite exponents in the denominator like this: \[\LARGE \frac{1}{x^n}=x^{-n}\]

OpenStudy (abmon98):

\[-\sin^2x/(\sin^2x)^2\] \[-2cosxsinx/\sin^4x\]

OpenStudy (anonymous):

or it will be better if u use Quotient rule

OpenStudy (abmon98):

If i used quotient rule \[1'*\sin^2x-(2cosxsinx)/\sin^4x\] @Yahoo!

OpenStudy (abmon98):

0*sin^2x-2cosxsinx/sin^4x

OpenStudy (anonymous):

yeah...but that sin^ 4 x is for all terms since here the 1st term is 0 u dont need to consider it

OpenStudy (abmon98):

okay thank you one more question if you don't mind

OpenStudy (anonymous):

Sure..

OpenStudy (abmon98):

\[\sin^2(2x)\] to find derivative do we use product rule?

OpenStudy (anonymous):

u can.... but u can do it with product rule also think of the way u diff sin^2x in above case

OpenStudy (anonymous):

there it was x here it is 2x

OpenStudy (abmon98):

so we differentiate sin^2(2x) sinx=cosx 2cosxsin2x

OpenStudy (anonymous):

nope differentiate sin^2(2x) = 2sin2x cos2x * 2 = 4 sin2x cos2x

OpenStudy (abmon98):

f(x)=sin^2x g(x)=2x use chain rule: f'(g(x))*g'(x)' 2cos*sin*2x*2

OpenStudy (abmon98):

e^(cos(x)^2) find derivative with respect to x